OpenStudy (mayaal):

GIVING MEDALS AND FAN! Find the vertex, focus, directrix, and focal width of the parabola. -1/12x^2= y

OpenStudy (mayaal):

@sangya21 @hartnn @aum @ganeshie8 @ZeldaOfHyrule

OpenStudy (anonymous):

-1/12 x^2=y or, x^2=-12y This is a parabola of the fourth standard. It can be compared to x^2=-4ay So, -4a=-12 a=3 Vertex C(0,0) Focus F(0,-a)= F(0,-3) Directrix: y+3=0

OpenStudy (mayaal):

so is the vertex -3?

OpenStudy (mayaal):

@Navket_Jha ?

OpenStudy (anonymous):

No, The vertex is a point. and according to this question its the origin